A baseball hit by a bat with an average force of 1000 N accelerates at 4.0 x 103 m/s2.

What is the ball’s mass?


1
Expert's answer
2021-10-31T18:10:07-0400
"F=ma,""m=\dfrac{F}{a}=\dfrac{1000\ N}{4.0\times10^3\ \dfrac{m}{s^2}}=0.25\ kg."